GANG CHENG

338 Jamesville Ave. 4D, Syracuse, NY 13210 (315)443-4889(o) (315)479-8767(h) gcheng@top.cis.syr.edu

\underline{GOALS}:

System Analyst, Application or System Programmer and Software Engineer, with emphasis on networking or SIMD/MIMD parallel systems.

\underline{EMPLOYMENT}:

January 1992 - Present : System Analyst, Research Assistant,

Northeast Parallel Architectures Center, Syracuse University, Syracuse, NY

August 1991 - January 1992 : Teaching Assistant,

School of Computer and Information Science, Syracuse University, Syracuse, NY

March 1991 - August 1991 : Programmer Analyst,

Information System Support Laboratory, Virginia Tech, VA.

June 1988 - January 1991 : Senior System Analyst,

Chengdu Institute of Computer Applications, Chinese Academy of Sciences.

August 1985 - May 1988 : Research and Teaching Assistant, Computer Science and Engineering Department, Huazhong University of Science and Technology, China.

June 1983 - July 1985 : Application Programmer, Chengdu Institute of Geography, Chinese Academy of Sciences.

\underline{EDUCATION}:

M.S., Computer and Information Science, Syracuse University, Syracuse, NY, January 1993 (GPA 3.9/4.0).

M.S., Computer Science and Engineering, Huazhong University of Science and Technology, China, May 1988 (GPA 9.7/10.0).

B.S., Computer Science and Engineering, Huazhong University of Science and Technology, China, May 1983 (GPA 9.6/10.0).

\underline{PROGRAMMING SKILLS}:

Connection Machine 2/5(CMfortran, C*, Prism, CMMD, CMX11), DECmpp-12000(MPfortran, MPL, MPPE), nCUBE2(NDB), TCP/IP, Express, PVM.

SUNstation, DECstation, IBM RS6000(UNIX, ULTRIX, AIX), VAX-11(VMS), IBM-PC(DOS), Macintosh.

C, C++, Fortran(Fortran77, Fortran90), Lisp, Prolog, Pascal, Cobol, Basic, Algol60, 8086/8088 and 8031/8051 Assembly, VHDL.

X Window(Xlib, X Toolkit, OSF/Motif Toolkit), AVS, SQL, dBASE-IV/Clipper, MS-Window, HALO, many other software packages on UNIX-based workstations.

\underline{HONORS}:

1. Graduate Scholarship, Graduate School, Syracuse University (1991-92).

2. Junior Visiting Research Fellow, Chinese Academy of Sciences (1989-90).

3. Member of Chinese Computer Federation.

\underline{MAJOR PROJECTS}:

1. \underline{Financial Modeling on Parallel Systems}. A set of stock option pricing models are implemented in Fortran90 and ported to massively parallel systems such as CM-2 and CM-5(in CMFortran), DECmpp-12000(in MPFortran and MPL), nCUBE2(in Fortran77+Express) and networked workstations(in Fortran77+Express).

2. \underline{An Interactive Visualization Environment on a Distributed High Performance System}. Written in C and AVS modules and using financial modeling as an example, application’s computing modules running on CM2/5, DECmpp-12000, SUN, DEC, IBM workstations are coupled, over a LAN(Ethernet), with network-based GUI running on a graphical workstation IBM RS6000 and managed under X-Window. This is a highly portable UNIX-based system.

3. \underline{A Vector-Based Spatial Database Management and Map Display System}. In Fortran77, Clipper and MS-Window and as a subsystem of the VirGIS (Virginia Geographic Information System) Project at Virginia Tech, this PC-based system allows the user to manage and generate high quality graphics of DLG maps in various ways.

4. \underline{”LISP+PROLOG” Interpreter-Compiler}. In C, the host of this system is a Lisp interpreter into which a Prolog compiler(as a knowledge base processor), an editor and a debugger are integrated. In this AI language, a user can program in either independent or mixed paradigms of Lisp and Prolog.

5. \underline{A Highway Toll-Collecting System}. In Intel 8031/8051 assembly and as a firmware in MCS-51(a microchip processor), it handles real-time processing of input signals from magnetic card reader/encoder which are data entry equipment at highway toll-station.

6. \underline{A Debris Flow Database System}. In dBASE-IV and C, this system was built for the management of debris flow disasters occurred in Southwest China since 1960’s.

7. \underline{Employee Management System}. In C/UNIX, this database management system provides the employer with usual management and query services for personnel administration and finance transaction.

SELECTED PUBLICATIONS:

1. A Large Scale Comparison of Option Pricing Models with Historical Market Data, in Proc. of the 4th Symposium on the Frontiers of Massively Parallel Computation, October, 1992, McLean, VA, pp. 420-426 (with K. Mills, M. Vinson).

2. An Interactive Visualization Environment for Financial Modeling on Heterogeneous Computing Systems, to appear in Proc. of the 6th SIAM Conference on Parallel Processing for Scientific Computing, March 1993, Norfolk, VA (with K. Mills, G. Fox).

3. Software Issues and Performance of a Stock Option Pricing Model on the Connection Machine-2 and DECmpp-12000, to appear in Proc. of the 5th Australian Supercomputing Conference, December, 1992 (with K. Mills, M. Vinson, S. Ranka, G. Fox).

4. A Functional + Logic Programming Language in Interpretation-Compilation Implementation, Lisp And Symbolic Computation: An International Journal, Vol. 5, No. 3, 1992, pp. 133-156 (with Y. Zhang).

5. The Implementation of An AI Language Unifying Functional and Logic Paradigms, M.S. Thesis, Computer Science and Engineering Department, Huazhong University of Science and Technology, China, 1988.

6. A Computer-Based Information System on Moving Elements of Debris Flow, in Proc. of the 1st Japan-China Symposium on Landslides and Debris Flows, 1989, Tokyo, Japan (with Z. Kang).

7. A Province-Wide Natural Disaster Precaution System: An Application of Geographic Information System to Natural Disaster Reduction, in Proc. of the 2nd International Workshop on GIS, Beijing, 1990 (with X. Huang).

8. A Highway Toll-Collecting System with Mag-Card Reader, Computer Applications, 1990 (with S. Huang).

\underline{REFERENCES}: Available Upon Request.